OpenStudy (anonymous):

Find the solution to the equation 4x - 8 = 20

OpenStudy (anonymous):

ok: 4x-8=20 +8 +8 4x=28 divide each side by 4

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

Do you understand how I got that?